Prove that the function f(x)= cosx-cos3x is periodc

1
Expert's answer
2022-04-14T04:28:25-0400

"f(x+2\pi)=cos(x+2\pi)+cos3(x+2\pi)=cos(x+2\pi)+cos(3x+6\pi)="

"=cosx+cos3x=f(x)."

Thus, f(x) is periodic with period "2\pi."